OpenStudy (anonymous):

What is the slope-intercept form of this equation? 6x + 4y = 8

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

awnser is b as in boy

OpenStudy (anonymous):

solve each equation by flinging square roots. if the equation has no real-# solution, or write no solution. #13. M^2=81

OpenStudy (anonymous):

what happen2 my stuff

OpenStudy (anonymous):

ok its solution the mwould have a solution becuase mwould half to be 9^2 to make 81. like 9times 9 equals 81 .so all your doing is just multiplying the same number .you just half to find it like 9*9=9^2. 9*9*9=9^3 9*9*9*9=9^4

OpenStudy (anonymous):

so square root is= ^2 on the end of this #. like 12^2 is 144